OpenStudy (anonymous):

What's the logarithmic equation 2=log(3x+4)

zepdrix (zepdrix):

Gimme one sec, I got somethin for ya

OpenStudy (anonymous):

All logs without a base written in are base 10, and logs are like the inverses of exponents

zepdrix (zepdrix):

Are you able to solve the problem once it's in exponential form?

OpenStudy (anonymous):

Thank you so much for that, @zepdrix! I usually have trouble when it comes to putting it in exponential form and then solving. However, I'll see what I can do. 10^2=3x+4 so 100=3x+4 subtract 4 from 100 and 4 96=3x divide by three on both sides. So, my answer would be 32?

OpenStudy (anonymous):

Could either of you maybe explain to me how we would convert 5=log3 (x^2+18) or e^x e^(x+1)=1 into an exponential equation and then solve?

zepdrix (zepdrix):

Try to draw the arrows like I was doing before! :O So the base 3 points to the 5, then the 5 points to the x^2+18. It's like a spiral that goes around. So the exponential form is: 3^5 = x^2+18

zepdrix (zepdrix):

So the other equation: e^x * e^(x+1) = 1 This one is uhhhhh a little tricky. So we'll apply a rule of exponents: When we multiply terms of similar bases, we ADD the exponents. So our left side becomes: e^(x+(x+1)) = 1 We can then rewrite 1 in terms of the exponential of base e. e^0 = 1 let's write our right side like this. e^(x+x+1) = e^0

zepdrix (zepdrix):

The bases are the same, which means the exponents have to also be equal. Set them equal to one another. x+x+1 = 0 Then it's pretty straight forward solving for x from there.

OpenStudy (anonymous):

Yes! Thank you so much for all your help!! Just to make sure, the final answer to that problem is x= -1/2, right? @zepdrix

OpenStudy (mathmale):

I'd like to present a different point of view: As samarasamara has pointed out, All logs without a base written in are base 10, and logs are like the inverses of exponents These two facts are instrumental in solving this present problem. Given 2 = log (3x+4) and the equation 10 = 10, Let 2 be the exponent of the left 10 and let log(3x+4) be the exponent of the right 10. 2 This reduces to 10 = 3x+4, or 100=3x+4, or 3x=96, or x=?? that's the solution.

zepdrix (zepdrix):

-1/2? Ya looks good!

OpenStudy (mathmale):

zepdrix: perhaps we'd better each check our respective "answers" in the original equation.

OpenStudy (mathmale):

log (3[32]+4) = log 100 = log 10^ 2 = 2. log (3[-1/2]+4) = log (1/2) = log 1 - log 2 = 0 - log 2, or approx. -0.693.